OpenStudy (anonymous):

I want help to solve square root of 2multiply by square root of 3

OpenStudy (anonymous):

As in: \[\sqrt{2}*\sqrt{3}\] You just multiply the insides so the answer is: \[\sqrt{6}\]

OpenStudy (anonymous):

with the ues of radical rule

OpenStudy (anonymous):

square root of 4m^21n^6. help me to solve

OpenStudy (raden):

like this : ?